OpenStudy (myname):

A block weighing 67.5 N rests on a plane inclined at 25.0° to the horizontal. A force F is applied to the object at 35.0° to the horizontal, pushing it upward on the plane. The coefficients of static and kinetic friction between the block and the plane are, respectively, 0.391 and 0.156.

OpenStudy (myname):

(b) What is the minimum value of F that will start the block moving up the plane?

OpenStudy (turingtest):

do you know how to find the normal force?

OpenStudy (myname):

yep thats easy

OpenStudy (myname):

N = m*g*Cos(25)

OpenStudy (turingtest):

so then that makes static friction easy to find... can you also find the force of gravity acting along the incline then?

OpenStudy (myname):

yep

OpenStudy (myname):

Wx = m*g*Sin(25)

OpenStudy (myname):

Yes i get the static friction

OpenStudy (turingtest):

okay, those are the two forces opposing motion of the block up the incline so the sum of those forces (or rather, infinitesimally over that) is the force required to overcome those two and push the block up the plane.

OpenStudy (myname):

So this is what i have: F -(Fs + Wx) = m*a

OpenStudy (myname):

here i dont have the acceleration

OpenStudy (myname):

i have 2 unknowns

OpenStudy (turingtest):

you just asked for F, nobody said anything about acceleration...

OpenStudy (myname):

sorry its Fr instead of just F so, Fr - (Fs+ Wx) = m*a

OpenStudy (myname):

if we had a then it would have been easier

OpenStudy (myname):

so where do i go from here

OpenStudy (turingtest):

whatever you call it, the point is that the block will start moving just past the point of maximum static friction, at which point the block is still not moving, so the forces sum to zero!

OpenStudy (myname):

I think i need to figure out another equation

OpenStudy (turingtest):

Fr - (Fs+ Wx)=? just before the block gets enough force to start moving, it's not moving. Therefore it's not accelerating. Therefore just befor the block moves the sum of the forces is...?

OpenStudy (myname):

oh so its 0

OpenStudy (myname):

Fr = (Fs + Wx)

OpenStudy (turingtest):

yep... at least that is right before it starts moving

OpenStudy (turingtest):

In a way I think the question is flawed because it would be infinitesimally little *more* force than that to actually get it to move, but I think we can overlook that here.

OpenStudy (myname):

and if it starts moving with constant velocity then --> Fr =(Fk+Wx) right

OpenStudy (myname):

Fk is the kinetic friction force

OpenStudy (turingtest):

...but it would have to accelerate from rest to get a velocity; it can't go from not moving to moving without accelerating, ya know what I mean? but again, I think they want you to see it the way you just did, so I would go with that. Fr =(Fk+Wx)

OpenStudy (myname):

you r right

OpenStudy (turingtest):

yeah, it's subtle if the forces are equal it could /keep/ moving with some constant velocity, but it can never change but the heck with that, the answer is Fr=(Fk+Wx)+dF (where dF is a differential force that gets it moving, so we can ignore it)

OpenStudy (myname):

Alright i will be back but in the time can you give me the right force for (b)

OpenStudy (turingtest):

just go with what you have...

OpenStudy (myname):

Cuz I got 53.24 N and it says its wrong

OpenStudy (myname):

this is what i had

OpenStudy (turingtest):

you can't use the kinetic force, it isn't moving yet! jeez this question has all kinds of subtleties it asks for the force to get it to *start* moving

OpenStudy (myname):

---> F*Cos(10) = Fs+Wx

OpenStudy (turingtest):

why cos(10) ?

OpenStudy (myname):

cuz the force is in an angle of 35 and the incline is in an angle of 25 with the horizontal

OpenStudy (myname):

U can post ur comments please . I willl be back in couple of minutes, alright.

OpenStudy (turingtest):

sure, and yeah, you are right about the cos(10) thing i missed that part

OpenStudy (myname):

I still cant get this right